OpenStudy (anonymous):

Linear equation A= x + y + z/3 solve for y

OpenStudy (anonymous):

I got the answer y = 3A-x-z but only cause i followed an example so im kind of confused.

OpenStudy (anonymous):

same here

OpenStudy (rizags):

thats not the right answer

OpenStudy (rizags):

first move all varibales to one side, leaving y alone

OpenStudy (anonymous):

On the answer sheet it said it was lol but my teacher said it could be more than one answer

OpenStudy (anonymous):

which is all ready done

OpenStudy (rizags):

Wait up, is it \[A= \frac{x+y+z}{3}\] or \[A= x+ y + \frac{z}{3}\]

OpenStudy (anonymous):

the first one

OpenStudy (rizags):

ahh, then, yes your answer is correct

OpenStudy (rizags):

first multiply both sides by 3 to get \[3A = x+y+z\]

OpenStudy (rizags):

then subtract x and z \[3A-x-z=y\]

OpenStudy (anonymous):

Thats the part i get confused where am i subtracting x and z?

OpenStudy (anonymous):

y=\[\frac{ a-z-x }{ 3 }\]

OpenStudy (anonymous):

no your not

OpenStudy (rizags):

from here: \[3A = x+y+z\] you subtract x and z

OpenStudy (anonymous):

no you divide by three to get y alone

OpenStudy (anonymous):

ohhh i see because youre trying to get the y alone

OpenStudy (anonymous):

yah

OpenStudy (anonymous):

and what ever is on the other side is y

OpenStudy (anonymous):

okay thank you both :)

OpenStudy (anonymous):

no problem